clouddiiskxyys:

All of the number sentences are true except. 3√1,331 = 11 3√125 = 5 3√64 = 8 3√1 = 1

NoodlesAndRiceYT:

first, let me ask you, what do you know about this problem?

surjithayer:

\[11^3=11 \times11\times 11=121 \times 11=1331\] \[5^3=5 \times 5 \times 5=25 \times 5=125\] \[8^3=8\times8\times8=64\times 8=512\] \[1^3=1\times 1\times 1=1\times 1=1\] can you guess the answer?